# On Minimal Achievable Quotas in Multiwinner Voting

### v1

@@ B.2 Complexity Results
3. Observation: Since for every candidate in $C_{X}$ there exists $a$ voters in $X$ that precisely approve that candidate, this implies that every committee $W$ that satisfies $\alpha$-JR has to contain all candidates $C_{X}$.
@@
With these observations, we can finally prove the equivalence: First, we show that if there exists a vertex cover for the original instance, then there exists a committee that satisfies $\alpha$-JR. So let $\{v^{VC}_{1},\dots v_{r}^{VC}\}$ be a vertex cover. We create the following committee
$W=C_{X}\cup\{c_{i}^{v_{i}^{VC}}\colon i\in[r]\}.$
@@
For the reverse direction, assume that there exists a committee $W$ that satisfies $\alpha$-JR. As argued above, it has to hold that $C_{X}\subseteq W\subseteq C_{X}\cup C_{Y}$. Further, it holds that at most $r$ voters $V^{\prime}\subseteq V$ approve any candidate in $W$ (at most one for every $c\in W\cap C_{y}$). We claim that $V^{\prime}$ forms a vertex cover. Assume for the sake of contradiction that this does not hold. Then, there exists an edge $e=\{u,v\}$ that is not covered by any vertex in $V^{\prime}$. Thus, neither $u$ nor $v$ approves any candidate in $W$ but both approve the common candidate $c_{e}^{E}$. Further, every voter in $F$ does not approve any candidate in $W$ and also approves $c_{e}^{E}$. Together they form an $\alpha$-JR violation contradicting the assumption that $W$ satisfies $\alpha$-JR. This concludes the proof for $\alpha$-JR.
In the following, we argue that the same reduction can also be used to show hardness for $\alpha$-EJR and $\alpha$-EJR+. For this, we analyze how many voters approve of the different candidates. It holds that $\lvert N_{c}\rvert=a$ for every $c\in C_{X}$ since $a$ voters in $X$ approve the candidate. For every candidate in $C_{E}$, it holds that all a2a-2 voters in $E$ and two voters in $V$ approve the candidate, so again, a total of $a$. For every candidate in $C_{y}$, $a$ voters in $Y$, and one voter in $V$ approve the candidate for a total of $a+1$ voters. Every 2-cohesive group must be of size at least
@@
First, assume that the ILP has a solution. Define $W=\{c\in C\colon x_{c}=1\}$ as our committee. We claim that $W$ does not have any $\alpha$-JR violation. Assume for the sake of contradiction, there exists one with witness $(N^{\prime},c^{\prime})$. This means all voters in $N^{\prime}$ approve $c^{\prime}$ and all voters in $N^{\prime}$ do not approve any candidate in $W$. Therefore, we can conclude from (1) that $x_{c}=0$ for all $c\in A_{i}$ with $i\in N^{\prime}$. Therefore, using (2), we get that $y_{i}=0$ for all $i\in N^{\prime}$. Finally, considering (3) for candidate $c^{\prime}$, we get that $\sum_{i\colon\,c^{\prime}\in A_{i}}(1-y_{i})\leq\lvert N^{\prime}\rvert$ and that $\lvert N^{\prime}\rvert\geq\lceil\alpha\cdot\frac{n}{k}\rceil$, which contradicts the assumption that the ILP had a solution for $\alpha$.
For the other direction, assume there is a committee $W$ that, for a given $\alpha$, satisfies $\alpha$-JR. We can define the following solution for the ILP. We set $x_{c}=1$ if and only if $c\in W$ and $y_{i}=1$ if and only if $W\cap A_{i}\neq\emptyset$ for all $c\in C,i\in N$. We claim that this is a valid solution for the ILP. First note that since $\lvert W\rvert=k$, condition (1) is satisfied. Consider now condition (2). If $y_{i}=0$, then the condition trivially holds, so assume that $y_{i}=1$. If $y_{i}=1$, then by definition there exists a candidate $c\in A_{i}\cap W$ and therefore $x_{c}=1$. So, finally assume for the sake of contradiction that (3) is violated for some $c$. Then, there exists a group $N^{\prime}$ such that $\lvert N^{\prime}\rvert\geq\lceil\alpha\cdot\frac{n}{k}\rceil$ and for every member $i\in N^{\prime}$, we have $y_{i}=0$ and $c\in A_{i}$. If $y_{i}=0$, this implies that $A_{i}\cap W=\emptyset$. But, then $(N^{\prime},c)$ are witnessing an $\alpha$-JR violation contradicting our assumption that $W$ satisfies $\alpha$-JR. This concludes the proof. ∎
@@
We give a reduction which can be used for basically all hardness results and counterexamples that already hold for some larger $\alpha$ (in this case $\alpha=1$).
We know that for smaller $\alpha$, all conditions get even harder to satisfy. To be able to use the same arguments as in the original instance ($\alpha=1$), we can rescale it accordingly.
Formally, let $\alpha=\frac{a}{b}$ with $a,b\in\mathbb{N}$. Given any instance $(N,C,k)$ for the Hare quota, we transform the instance to the following instance $(N_{\alpha},C,k)$ such that $N_{\alpha}=\bigcup_{i=1}^{a}N^{(i)}\cup Z$, where each $N^{(i)}$ is an exact copy of the voters in $N$ meaning that the voters approve precisely the same candidates in $C$, and $Z=\{z_{1},\dots z_{(b-a)\cdot\lvert N\rvert}\}$ are voters that do not approve any candidate in $C$.
@@
For the reverse direction, let $(N^{\prime}_{\alpha},C^{\prime},\ell)$ be an EJR violation. Thus, it holds that $\lvert N^{\prime}_{\alpha}\rvert\geq\alpha\cdot\ell\cdot\frac{n_{\alpha}}{k}=\ell\cdot\frac{n\cdot a\cdot b}{k\cdot b}=a\cdot\ell\cdot\frac{n}{k}$. We now define the set
@@
First, observe that $\lvert N^{\prime}\rvert\geq\frac{\lvert N^{\prime}_{\alpha}\rvert}{a}=\ell\cdot\frac{n}{k}$. Furthermore, no voter in $N^{\prime}$ approves more than $\ell-1$ voters in $W$. Thus, $(N^{\prime},C,\ell)$ witnesses an EJR violation for the original instance. This concludes the proof. ∎
@@
The proof resembles the one for Proposition˜5.1. We compute for every candidate $c\notin W$ and $\ell\in\{1,\dots,k\}$ the set of all voters that approve the candidate $c$ but approve at most $\ell-1$ candidates in $W$. Formally,
@@
Observe that this group of voters $N^{\prime}_{c,\ell}$ forms an $\alpha$-EJR violation for $\alpha=\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}$ since $\lvert N^{\prime}_{c,\ell}\rvert\geq\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}\cdot\ell\cdot\frac{n}{k}=\alpha\cdot\ell\cdot\frac{n}{k}$.
Finally, let $\alpha=\max_{c\notin W,\ell\in\{1,\dots,k\}}\;\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}$. Therefore, $\alpha\leq\alpha_{\text{EJR+}}$.
It remains to show that $\alpha\geq\alpha_{\text{EJR+}}$. By definition, there exists some $\alpha_{\text{EJR+}}$-EJR+ violation $(N^{\prime},c,\ell)$ for $N^{\prime}\subseteq N$ and $c\in C$. Thus, it holds that $\lvert N^{\prime}\rvert\geq\alpha_{\text{EJR+}}\cdot\ell\cdot\frac{n}{k}$. Further, since the voters in $N^{\prime}$ all approve $c$ but at most $\ell-1$ candidates in $W$, we get that $\lvert N^{\prime}\rvert\leq\lvert\{i\in N_{c}\colon\lvert A_{i}\cap W\rvert<\ell\}\rvert\leq\lvert N^{\prime}_{c,\ell}\rvert$. Consequently, $\lvert N^{\prime}_{c,\ell}\rvert\geq\alpha_{\text{EJR+}}\cdot\ell\cdot\frac{n}{k}$ which implies that $\alpha_{\text{EJR+}}\leq\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}\leq\alpha$. This concludes the proof. ∎
@@ B.3 Restricted Domains
Now, to find the largest violation, it is sufficient to consider every possible interval of which $O(n^{2})$ many exist.
Trivially, every violation we find with this procedure is an $\alpha$-EJR violation. In the following, we also prove that if there exists an $\alpha$-EJR violation $(N^{\prime},C^{\prime},\ell)$ then the algorithm above finds an $\alpha$-EJR violation $(N^{\prime\prime},C^{\prime\prime},\ell^{\prime})$ with $\ell^{\prime}\geq\ell$, and $N^{\prime}\subseteq N^{\prime\prime}$.
So let $N^{\prime}$ be the set of voters witnessing the violation, and let $v_{i}$ be the first voter, as well as $v_{j}$ be the last voter in $N^{\prime}$, according to the VI order. When the algorithm considered the interval $(i,j)$, there were at least $\lvert N^{\prime}\rvert\geq\frac{n}{k}\cdot\ell\cdot\alpha$ many voters approving at most $\ell-1$ many candidates in $W$. Further, each of them approves at least $\ell$ common candidates $C^{\prime}$ with $C^{\prime}\subseteq T(i,j)$. Thus, it holds that the algorithm finds a violation of size at least $\ell$ for this interval $(i,j)$. This concludes the proof for EJR.
For JR, the same arguments work if we only consider $1$-cohesive groups. ∎
@@
To prove the theorem, we use Algorithm˜1. Observe that the algorithm itself does not necessarily return a subset of size $k$. The algorithm returns the smallest subset $W$ that does not contain any $\alpha$-JR violation. Lemma˜B.2 proves this statement. Thus, we can execute Algorithm˜1 and a committee satisfying $\alpha$-JR exists if and only if the algorithm returns a subset with at most $k$ candidates. ∎
@@
We prove this statement by induction over the voters according to the VI order. We say a subset of candidates is feasible if it satisfies $\alpha$-JR and optimal if it is a feasible subset with minimal cardinality. Observe that the algorithm returns a feasible subset of candidates, so there always exists at least one optimal feasible subset of candidates. Let $C_{j}$ be the set of candidates selected up to timestep $j$ from our algorithm.
Claim. For every $j\in\{1,\dots,n\}$, there exists an optimal subset of candidates $W^{(j)}$ such that every candidate that the greedy algorithm selected at some index $\leq j$ belongs to $W^{(j)}$, i.e., $C_{i}\subseteq W^{(j)}$.
Base: $j=1$. If $C_{i}=\emptyset$, we are done, so assume that the set is non-empty. This means voter $v_{1}$ causes an $\alpha$-JR violation by herself, and every feasible subset of candidates must select a candidate that is approved by voter $v_{1}$. Observe that the candidate $c$ selected by Algorithm˜1 is defined such that $r_{c^{\prime}}\sqsubseteq r_{c}$ for all $c^{\prime}\in A_{v_{1}}$. Since we are in the VI domain, this implies that $N_{c^{\prime}}\subseteq N_{c}$. Therefore, if an optimal subset does not contain $c$ it has to contain such a $c^{\prime}$, and we can swap them while upholding $\alpha$-JR. This proves the induction base.
Step: Let the claim be satisfied for some $j-1\geq 0$. If no candidate is selected in iteration $j$, we have $C_{j-1}=C_{j}$ and the claim follows from the induction hypothesis since $C_{j-1}\subseteq W^{(j-1)}$. So assume a candidate $c_{j}$ is selected in iteration $j$. Since there is an $\alpha$-JR violation caused by a set of voters $N^{\prime\prime}$ with $v\sqsubseteq v_{j}$ for every $v\in N^{\prime\prime}$ and containing voter $v_{j}$, $W^{(j-1)}\setminus C_{j-1}$ must contain a candidate $c^{\prime\prime}$ that is approved by some voter $v\in N^{\prime\prime}$ and thus $v\sqsubseteq v_{j}$.
@@
It remains to consider the case that $v^{\prime\prime}\sqsubset v_{j}$. Now, since by definition $C_{j-1}$ does not contain any $\alpha$-JR violation for $\{v_{1},\dots,v_{i-1}\}$, and $C_{j-1}\subseteq W^{(j)}$ there has to exist a voter $v^{*}\in N^{\prime}$ such that $v_{j}\sqsubseteq v^{*}$. But then $v^{\prime\prime}\sqsubset v_{j}\sqsubseteq v^{*}$ which implies that $c\in A_{v_{j}}$. Similar to the first case, this implies that $r_{c}\sqsubseteq r_{c_{j}}$ and thus it holds that every voter $v^{*}$ with $v_{j}\sqsubseteq v^{*}$ that approves $c$ also approves $c_{j}$. But this is a contradiction to $v^{*}\in N^{\prime}$.
Consequently, $W^{(n)}$ is minimal and feasible and corresponds to the output of our algorithm. Therefore, there cannot be any feasible committee $W^{\prime}$ with $\lvert W^{\prime}\rvert<\lvert W\rvert$. ∎
@@
We can use a similar proof as for the VI domain, but instead of considering every interval of voters, we consider every interval of candidates.
More precisely, denote by $N_{\ell}^{i,j}=\{v\in N:\{c_{i},\dots,c_{j}\}\subseteq A_{v},\lvert A_{v}\cap W\rvert$$\ell\}$ for $1\leq\ell\leq k$ the voters that approve every candidate in the interval but at most $\ell-1$ many candidates in $W$.
Then, denote by $\ell_{i,j}=\max_{\ell}(\min((j-i+1,\lfloor\lvert N_{\ell}^{i,j}\rvert\cdot\frac{k}{n}\cdot\alpha\rfloor)\geq\ell)$. Observe that if there does not exist any cohesive group, the value is equal to 0. We claim that with this procedure, we can find an EJR violation if one exists.
To prove this, let $(N^{\prime},C^{\prime},\ell)$ be any EJR violation for the committee $W$. Further, let $c_{i}$ and $c_{j}$ be the first and last candidates in $C^{\prime}$ according to the CI order. By definition, every voter in $N^{\prime}$ approves at most $\ell-1$ many candidates in $W$ and $\lvert C^{\prime}\rvert\geq\ell$. Thus, also $(N^{\prime},\{c_{i},\dots,c_{j}\},\ell)$ is a witness of an EJR violation. When the algorithm considers the interval $i,j$, it therefore finds a violation $(N^{\prime\prime},\{c_{i},\dots,c_{j}\},\ell)$ with $N^{\prime}\subseteq N^{\prime\prime}$. This concludes the proof. ∎
@@
We claim that Algorithm˜2 returns a subset of size at most $k$ if and only if there exists a committee of size $k$ that satisfies $\alpha$-JR. First, observe that Algorithm˜2 always outputs a subset of candidates that satisfies $\alpha$-JR by construction. So if the output committee is of size at most $k$, there trivially also exists a committee satisfying $\alpha$-JR. It remains to show that if there exists a committee of size $k$, then the algorithm finds one as well.
@@
Claim. For a given $j$ exists a minimal committee $W^{(j)}$ that satisfies $\alpha$-JR and $C_{j}\subseteq W^{(j)}$.
@@
Base: $j=1$. As discussed above, the subset of candidates returned by the algorithm satisfies $\alpha$-EJR. Therefore, there exists a subset of candidates that satisfies $\alpha$-JR and therefore also one of minimal cardinality. Now, if $c_{1}$ is selected by our algorithm, then $c_{1}$ must also be in every optimal committee since the set of candidates $\{c_{2},\dots,c_{m}\}$ and therefore all possible subsets of size $k$ of this set would result in an $\alpha$-JR violation. Therefore, $C_{1}\subseteq W^{(1)}$.

### v2

@@ B.2 Complexity Results
3. Observation: Since for every candidate in $C_{X}$ there exist $a$ voters in $X$ that precisely approve that candidate, this implies that every committee $W$ that satisfies $\alpha$-JR has to contain all candidates $C_{X}$.
@@
With these observations, we can finally prove the equivalence: First, we show that if there exists a vertex cover for the original instance, then there exists a committee that satisfies $\alpha$-JR. So let $\{v^{\text{VC}}_{1},\dots v_{r}^{\text{VC}}\}$ be a vertex cover. We create the following committee
$W=C_{X}\cup\{c_{i}^{v_{i}^{\text{VC}}}\colon i\in[r]\}.$
@@
For the reverse direction, assume that there exists a committee $W$ that satisfies $\alpha$-JR. As argued above, it has to hold that $C_{X}\subseteq W\subseteq C_{X}\cup C_{Y}$. Further, it holds that at most $r$ voters $V^{\prime}\subseteq V$ approve any candidate in $W$ (at most one for every $c\in W\cap C_{y}$). We claim that $V^{\prime}$ forms a vertex cover. Assume for the sake of contradiction that this does not hold. Then, there exists an edge $e=\{u,v\}$ that is not covered by any vertex in $V^{\prime}$. Thus, neither $u$ nor $v$ approves any candidate in $W$, but both approve the common candidate $c_{e}^{E}$. Further, every voter in $F$ does not approve any candidate in $W$ and also approves $c_{e}^{E}$. Together they form an $\alpha$-JR violation contradicting the assumption that $W$ satisfies $\alpha$-JR. This concludes the proof for $\alpha$-JR.
In the following, we argue that the same reduction can also be used to show hardness for $\alpha$-EJR and $\alpha$-EJR+. To this end, we analyze the number of voters who approve of each candidate. It holds that $\lvert N_{c}\rvert=a$ for every $c\in C_{X}$ since $a$ voters in $X$ approve the candidate. For every candidate in $C_{E}$, all a2a-2 voters in $E$ and two voters in $V$ approve the candidate, so again, a total of $a$. For every candidate in $C_{y}$, $a$ voters in $Y$, and $1$ voter in $V$ approve the candidate for a total of $a+1$ voters. Every 2-cohesive group must be of size at least
@@
First, assume that the ILP has a solution. Let this be $x\in\{0,1\}^{m}$ and $y\in\{0,1\}^{n}$. Define $W=\{c\in C\colon x_{c}=1\}$ as our committee. We claim that $W$ does not have any $\alpha$-JR violation. Assume for the sake of contradiction, there exists one with witness $(N^{\prime},c^{\prime})$. This means all voters in $N^{\prime}$ approve $c^{\prime}$ and all voters in $N^{\prime}$ do not approve any candidate in $W$. Therefore, we can conclude from (1) that $x_{c}=0$ for all $c\in A_{v}$ with $v\in N^{\prime}$. Therefore, using (2), we get that $y_{v}=0$ for all $v\in N^{\prime}$. Finally, considering (3) for candidate $c^{\prime}$, we get that $\sum_{v\colon\,c^{\prime}\in A_{v}}(1-y_{v})\leq\lvert N^{\prime}\rvert$ and that $\lvert N^{\prime}\rvert\geq\lceil\alpha\cdot\frac{n}{k}\rceil$, which contradicts that $x$ and $y$ being a solution for the ILP.
For the other direction, assume there is a committee $W$ that, for a given $\alpha$, satisfies $\alpha$-JR. We can define the following solution for the ILP. We set $x_{c}=1$ if and only if $c\in W$ and $y_{v}=1$ if and only if $W\cap A_{v}\neq\emptyset$ for all $c\in C,v\in N$. We claim that this is a valid solution for the ILP. First note that since $\lvert W\rvert=k$, condition (1) is satisfied. Consider now condition (2). If $y_{v}=0$, then the condition trivially holds, so assume that $y_{v}=1$. If $y_{v}=1$, then by definition there exists a candidate $c\in A_{v}\cap W$ and therefore $x_{c}=1$. So, finally, assume for the sake of contradiction that (3) is violated for some $c$. Then, there exists a group $N^{\prime}$ such that $\lvert N^{\prime}\rvert\geq\lceil\alpha\cdot\frac{n}{k}\rceil$ and for every member $v\in N^{\prime}$, we have $y_{v}=0$ and $c\in A_{v}$. If $y_{v}=0$, this implies that $A_{v}\cap W=\emptyset$. But, then $(N^{\prime},c)$ are witnessing an $\alpha$-JR violation contradicting our assumption that $W$ satisfies $\alpha$-JR. This concludes the proof. ∎
@@
We give a reduction that applies to essentially all hardness results and counterexamples that already hold for some larger $\alpha$ (in this case, $\alpha=1$).
We know that for smaller $\alpha$, all conditions get even harder to satisfy. To use the same arguments as in the original instance ($\alpha=1$), we can rescale accordingly.
Formally, let $\alpha=\frac{a}{b}$ with $a,b\in\mathbb{N}$ (possible by Lemma˜4.11). Given any instance $(N,C,k)$ for the Hare quota, we transform the instance to the following instance $(N_{\alpha},C,k)$ such that $N_{\alpha}=\bigcup_{i=1}^{a}N^{(i)}\cup Z$, where each $N^{(i)}$ is an exact copy of the voters in $N$ meaning that the voters approve precisely the same candidates in $C$, and $Z=\{z_{1},\dots z_{(b-a)\cdot\lvert N\rvert}\}$ are voters that do not approve any candidate in $C$.
@@
The proof resembles the one from Proposition˜5.1. We compute for every candidate $c\notin W$ and $\ell\in\{1,\dots,k\}$ the set of all voters that approve the candidate $c$ but approve at most $\ell-1$ candidates in $W$. Formally,
@@
Observe that this group of voters $N^{\prime}_{c,\ell}$ forms an $\alpha$-EJR+ violation for $\alpha=\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}$ since $\lvert N^{\prime}_{c,\ell}\rvert=\alpha\cdot\ell\cdot\frac{n}{k}$.
Finally, let $\alpha=\max_{c\notin W,\ell\in\{1,\dots,k\}}\;\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}$. Therefore, $\alpha\leq\alpha_{\text{EJR+}}^{W}$.
It remains to show that $\alpha\geq\alpha_{\text{EJR+}}^{W}(I)$. By definition, there exists some $\alpha_{\text{EJR+}}^{W}(I)$-EJR+ violation $(N^{\prime},c,\ell)$ for $N^{\prime}\subseteq N$ and $c\in C$. Thus, it holds that $\lvert N^{\prime}\rvert\geq\alpha_{\text{EJR+}}^{W}(I)\cdot\ell\cdot\frac{n}{k}$. Further, since the voters in $N^{\prime}$ all approve $c$ but at most $\ell-1$ candidates in $W$, we get that $\lvert N^{\prime}\rvert\leq\lvert\{i\in N_{c}\colon\lvert A_{i}\cap W\rvert<\ell\}\rvert\leq\lvert N^{\prime}_{c,\ell}\rvert$. Consequently, $\lvert N^{\prime}_{c,\ell}\rvert\geq\alpha_{\text{EJR+}}^{W}(I)\cdot\ell\cdot\frac{n}{k}$ which implies that $\alpha_{\text{EJR+}}^{W}(I)\leq\lvert N^{\prime}_{c,\ell}\rvert\cdot\frac{k}{\ell n}\leq\alpha$. This concludes the proof. The runtime is $O(mnk)$, since for every $c\notin W$ and $l\leq k$, we need to determine the number of voters who approve the candidate $c$. ∎
@@ B.3 Restricted Domains
See 6.2
Proof.
We first show that $\alpha$-EJR+ and $\alpha$-EJR are equivalent for party-list profiles. The first direction follows from Proposition˜4.7, so we focus on the other one: Let $(N^{\prime},c,\ell)$ be an EJR+ violation, i.e., $\lvert N^{\prime}\rvert\geq\alpha\ell\frac{n}{k}$, $c\in(C\setminus W)\cap(\bigcap_{v\in N^{\prime}}A_{v})$ and for all $v\in N^{\prime}$, we have $\lvert W\cap A_{v}\rvert$$\ell$ and $\lvert\bigcap_{v\in N^{\prime}}A_{v}\rvert\geq\ell$. This follows from the fact that all voters in $N^{\prime}$ approve the same set of candidates; therefore, this also forms an $\alpha$-EJR violation. In the following, we will demonstrate the theorem for $\alpha$-EJR, which, as shown previously, also holds for $\alpha$-EJR+.
$\arg\max_{p\in P}\frac{s_{p}}{w_{p,i}+1}$
where $w_{p,i}$ is the number of candidates approved by party $p$ in the committee $W$ at timestep $i$. The intuition behind the algorithm is the following: At every timestep, check which party causes the largest violation.
For the party $p$ not to get $(\alpha,w_{p,i}+1)$-cohesive, it has to hold that $\frac{n}{k}\cdot\alpha\cdot(w_{p,i}+1)>s_{p}$. Solving for $\alpha$, we get that $\alpha>\frac{s_{p}}{w_{p,i}+1}\cdot\frac{k}{n}$ where $\frac{k}{n}$ is a constant factor for all parties and can thus be ignored.
This contradicts our assumption that $\alpha^{\prime}<\alpha$.
In order to efficiently execute the algorithm we use a heap to store the parties. Then, in every one of the $k$ rounds, we need to compute the current value $\frac{s_{p}}{w_{p,i}+1}$ for every party $p\in P$. This can be implemented using a max-heap, yielding a runtime of $)O(\lvert P\rvert+k\log(\lvert P\rvert$)).
For $\alpha$-JR we can find an optimal committee by making out the $k$ parties with the largest support and selecting a candidate from each of them. In order to find $\alpha_{\text{JR}}^{*}(I)$ we consider the size of the $k+1$ largest party $P_{k+1}$ and compute $\alpha_{\text{JR}}^{*}(I)=\lvert P_{k+1}\rvert\cdot\frac{k}{n}$. Both can be done in linear time (Blum et al., 1973). ∎
@@
For each $1\leq\ell\leq\lvert T(i,j)\rvert$, we compute the number of voters that approve at most $\ell-1$ many candidates in $W$. If this group is large enough to witness an $\alpha$-EJR violation, we know that $W$ does not satisfy $\alpha$-EJR since every voter in this group approves every candidate in $T(i,j)$. Finding a violation with this procedure, therefore, takes $O(nk)$ time for a given interval $(i,j)$. It is sufficient to consider every possible interval of which $O(n^{2})$ many exist. Thus, the algorithm’s runtime is in $O(n^{3}k)$.
Trivially, every violation we find with this procedure is an $\alpha$-EJR violation. In the following, we also prove that if there exists an $\alpha$-EJR violation $(N^{\prime},C^{\prime},\ell)$ then the algorithm above finds an $\alpha$-EJR violation $(N^{\prime\prime},C^{\prime\prime},\ell)$ with $N^{\prime}\subseteq N^{\prime\prime}$.
So let $N^{\prime}$ be the set of voters witnessing the violation, and let $v_{i}$ be the first voter, as well as $v_{j}$ be the last voter in $N^{\prime}$, according to the VI order. When the algorithm considered the interval $(i,j)$, there were at least $\lvert N^{\prime}\rvert\geq\frac{n}{k}\cdot\ell\cdot\alpha$ many voters approving at most $\ell-1$ many candidates in $W$. Further, each of them approves at least $\ell$ common candidates $C^{\prime}$ with $C^{\prime}\subseteq T(i,j)$. Thus, the algorithm finds a violation in the interval $(i,j)$. This concludes the proof for EJR. ∎
@@
To prove the theorem, we use Algorithm˜1. Observe that the algorithm itself does not necessarily return a subset of size $k$. By Lemma˜B.2, it holds that the algorithm returns the smallest subset $W$ that does not contain any $\alpha$-JR violation. Thus, we can execute Algorithm˜1 and a committee satisfying $\alpha$-JR exists if and only if the algorithm returns a subset with at most $k$ candidates.
The runtime of the algorithm is $O(n^{2}m\log n)$ as in every one of the $n$ timesteps, we need to check if $\alpha$-JR is violated for a given committee $W$ and $\alpha$-value, which can be done in $O(nm)$ (shown in Proposition˜5.1). Using binary search over the distinct $\alpha$-values, of which at most $n$ exist due to Lemma˜4.10 requires an additional $\log n$ factor. ∎
@@
We prove this statement by induction over the voters, following the VI order. We say a subset of candidates is feasible if it satisfies $\alpha$-JR. Further, we say that a committee is optimal if it is feasible and has minimal cardinality. Observe that the algorithm returns a feasible subset of candidates, so there always exists at least one optimal feasible subset. Let $C_{j}$ be the set of candidates selected up to time step $j$ by our algorithm.
Claim. For every $j\in\{1,\dots,n\}$, there exists an optimal subset of candidates $W^{(j)}$ such that every candidate that the greedy algorithm selected at some index $\leq j$ belongs to $W^{(j)}$, i.e., $C_{j}\subseteq W^{(j)}$.
Base: $j=1$. If $C_{j}=\emptyset$, we are done, so assume that the set is non-empty. This means voter $v_{1}$ causes an $\alpha$-JR violation by herself, and every feasible subset of candidates must select a candidate that is approved by voter $v_{1}$. Observe that the candidate $c$ selected by Algorithm˜1 is defined such that $r_{c^{\prime}}\sqsubseteq r_{c}$ for all $c^{\prime}\in A_{v_{1}}$. Since we are in the VI domain, this implies that $N_{c^{\prime}}\subseteq N_{c}$. Therefore, if an optimal subset does not contain $c$, it has to contain such a $c^{\prime}$, and we can swap them while upholding $\alpha$-JR. This proves the induction base.
Step: Let the claim be satisfied for some $j-1\geq 0$. If no candidate is selected in iteration $j$, we have $C_{j-1}=C_{j}$ and the claim follows from the induction hypothesis since $C_{j-1}\subseteq W^{(j-1)}$ and setting $W^{(j)}=W^{(j-1)}$. So assume a candidate $c_{j}$ is selected in iteration $j$. Further, if $c_{j}\in W^{(j-1)}$ then $C_{j}\subseteq W^{(j-1)}$ which proves the claim. So in the following assume that $c_{j}\not\in W^{(j-1)}$.
Since there is an $\alpha$-JR violation caused by a set of voters $N^{\prime\prime}$ with $v\sqsubseteq v_{j}$ for every $v\in N^{\prime\prime}$ and containing voter $v_{j}$, $W^{(j-1)}\setminus C_{j-1}$ must contain a candidate $c^{\prime\prime}$ that is approved by some voter $v\in N^{\prime\prime}$ and thus $v\sqsubseteq v_{j}$.
@@
It remains to consider the case that $v^{\prime\prime}\sqsubset v_{j}$. Now, since by definition $C_{j-1}$ does not contain any $\alpha$-JR violation for $\{v_{1},\dots,v_{j-1}\}$, and $C_{j-1}\subseteq W^{(j)}$ there has to exist a voter $v^{*}\in N^{\prime}$ such that $v_{j}\sqsubseteq v^{*}$. But then $v^{\prime\prime}\sqsubset v_{j}\sqsubseteq v^{*}$ which implies that $c\in A_{v_{j}}$. As in the first case, this implies that $r_{c}\sqsubseteq r_{c_{j}}$, and thus every voter $v^{*}$ with $v_{j}\sqsubseteq v^{*}$ who approves $c$ also approves $c_{j}$. But this is a contradiction to $v^{*}\in N^{\prime}$.
Consequently, $C_{n}$ is both minimal and feasible, corresponding to the output of our algorithm. Therefore, there cannot be any feasible committee $W^{\prime}$ with $\lvert W^{\prime}\rvert<\lvert W\rvert$. ∎
@@
We can use a similar proof as for the VI domain, but instead of considering every interval of voters, we consider every interval of candidates. More precisely, denote by $N_{\ell}^{i,j}=\{v\in N:\{c_{i},\dots,c_{j}\}\subseteq A_{v},\lvert A_{v}\cap W\rvert$$\ell\}$ for $1\leq\ell\leq j-i+1$ the voters that approve every candidate in the interval but at most $\ell-1$ many candidates in $W$. If any such group $N_{\ell}^{i,j}$ is larger than $\alpha\ell\frac{n}{k}$ this group trivially witnesses an $\alpha$-EJR violation $(N_{\ell}^{i,j},\{c_{i},\dots,c_{j}\},\ell)$. Since we can compute $N_{\ell}^{i,j}$ in $O(n)$ and we consider at most $k$ values for $\ell$ for at most $O(m^{2})$ many pairs $i,j$ the runtime of the algorithm is at most $O(m^{2}kn)$. We claim that this procedure also always finds an $\alpha$-EJR violation whenever one exists.
To prove this, let $(N^{\prime},C^{\prime},\ell)$ be any $\alpha$-EJR violation for the committee $W$. Further, let $c_{i}$ and $c_{j}$ be the first and last candidates in $C^{\prime}$ according to the CI order. By definition, every voter in $N^{\prime}$ approves at most $\ell-1$ many candidates in $W$ and $\lvert C^{\prime}\rvert\geq\ell$. Thus, also $(N^{\prime},\{c_{i},\dots,c_{j}\},\ell)$ is a witness of an EJR violation. When the algorithm considers the interval $i,j$, it therefore finds a violation $(N^{\prime\prime},\{c_{i},\dots,c_{j}\},\ell)$ with $N^{\prime}\subseteq N^{\prime\prime}$. This concludes the proof. ∎
@@
We claim that Algorithm˜2 returns a subset of size at most $k$ if and only if there exists a committee of size $k$ that satisfies $\alpha$-JR. First, observe that Algorithm˜2 always outputs a subset of candidates that satisfies $\alpha$-JR by construction. So if the output committee is of size at most $k$, there trivially also exists a committee satisfying $\alpha$-JR. It remains to show that if a committee of size $k$ exists, then the algorithm also finds one.
@@
Claim. For a given $j$, there exists a minimal committee $W^{(j)}$ that satisfies $\alpha$-JR and $C_{j}\subseteq W^{(j)}$.
@@
Base: $j=1$. As discussed above, the subset of candidates returned by the algorithm satisfies $\alpha$-JR. Therefore, there exists a subset of candidates that satisfies $\alpha$-JR and thus also one of minimal cardinality. Now, if $c_{1}$ is selected by our algorithm, then $c_{1}$ must also be in every optimal committee since the set of candidates $\{c_{2},\dots,c_{m}\}$ and therefore all possible subsets of size $k$ of this set would result in an $\alpha$-JR violation. Therefore, $C_{1}\subseteq W^{(1)}$ for some optimal committee $W^{(1)}$.